OpenStudy (anonymous):

Solve by completing the square and then use the square root method to solve. x2 - 4x - 32 = 0

OpenStudy (anonymous):

(x - 2)^2 - 4 - 32 = 0 (x-2)^2 = 36 x-2 = 6 or -6 solution is 8, -4

OpenStudy (campbell_st):

rewrite the equation x^2 - 4x = 32 then add (4/2)^2 to both sides x^2 - 4x + 4 = 32 + 4 factorise (x-2)^2 = 36 \[x - 2 = \sqrt{36}\] \[x - 2 = \pm 6\] \[x = 2 \pm 6\] x = 2 + 6 x = 2 - 6 = 8 = -4